# Emergent Hydrodynamic Mode on SU(2) Plaquette Chains and Quantum Simulation

### v1

@@ Abstract
We search for emergent hydrodynamic modes in real-time Hamiltonian dynamics of $2+1$-dimensional SU(2) lattice gauge theory on a quasi one dimensional plaquette chain, by numerically computing symmetric correlation functions of energy densities on lattice sizes of about $20$ with the local Hilbert space truncated at $j_{\rm max}=\frac{1}{2}$. Due to the Umklapp processes, we only find a mode for energy diffusion. The symmetric correlator exhibits transport peak near zero frequency with a width proportional to momentum squared at small momentum, when the system is fully quantum ergodic, as indicated by the eigenenergy level statistics. This transport peak leads to a power-law $t^{-\frac{1}{2}}$ decay of the symmetric correlator at late time, also known as the long-time tail, as well as diffusion-like spreading in position space. We also introduce a quantum algorithm for computing the symmetric correlator on a quantum computer and find it gives results consistent with exact diagonalization when tested on the IBM emulator. Finally we discuss the future prospect of searching for the sound modes.
@@ I Introduction
The values of shear viscosity and other transport coefficients of QCD at varying temperatures serve as important inputs for an array of phenomenological studies in heavy ion collisions. However, calculating them accurately in QCD is challenging due to nonperturbative effects and the sign problem in Euclidean lattice QCD approach [13], although progress has been made [14, 15, 16, 17, 18, 19, 20, 21, 22]. This challenge together with the rapid development of quantum computing technology motivated recent studies on using the Hamiltonian formulation of lattice gauge theory to calculate the shear viscosity [23, 24]. The Hamiltonian lattice method is limited to only small systems at the moment [24], due to the limitation of current classical computers and quantum devices, which can lead to large systematic uncertainty. In particular, the lattice size may be too small to support the full development of long-wavelength hydrodynamic behavior. In this paper, we aim at answering this question and demonstrating one case where hydrodynamic behavior emerges in the Hamiltonian dynamics on the lattice. Because of the real-time nature, such a demonstration is difficult in the Euclidean lattice setup.
The main object that we will calculate in the Hamiltonian lattice approach is the real-time symmetric correlation function of stress-energy tensors, which is expected to evolve hydrodynamically at late time. Our results will show that a diffusive mode emerges in the SU(2) Hamiltonian dynamics on a lattice of size $\sim 20$. Due to the Umklapp effect, sound modes are not observed at the couplings we study with the local Hilbert space truncation.
@@ II.1 Hydrodynamics in the continuum
$\delta T^{\mu\nu}$ $=\begin{pmatrix}\delta\varepsilon_{0}&-(\varepsilon_{0}+P_{0})\delta u^{x}\\ -(\varepsilon_{0}+P_{0})\delta u^{x}&\delta P-\zeta\partial_{x}\delta u^{x}\end{pmatrix}\,.$ (4)
@@
$\partial_{t}\delta\varepsilon-\partial_{x}g^{x}=0\,,$
$\partial_{t}g^{x}-c_{s}^{2}\partial_{x}\delta\varepsilon+\gamma_{\zeta}\partial_{x}^{2}g^{x}=0\,,$ (6)
@@
$\partial_{t}^{2}g^{x}-c_{s}^{2}\partial_{x}^{2}g^{x}+\gamma_{\zeta}\partial_{t}\partial_{x}^{2}g^{x}=0\,.$ (7)
@@
$\omega_{s\pm}=\pm\sqrt{c_{s}^{2}k^{2}-\frac{\gamma_{\zeta}^{2}k^{4}}{4}}+i\frac{\gamma_{\zeta}k^{2}}{2}\,.$ (9)
@@ II.3 Real-time symmetric correlator
where $\rho_{T}\equiv\frac{e^{-\beta H}}{Z}$ with $\beta=1/T$ denotes the thermal density matrix, $t_{0}$ and $x_{0}$ are some arbitrary reference points that can be set to zeros. The symmetric correlator characterizes the fluctuations in the system at thermal equilibrium.
@@ III Hamiltonian of SU(2) Plaquette Chains
We will compute the symmetric correlator of the energy density for a specific lattice gauge theory, the SU(2) pure gauge theory in $2+1$ dimensions discretized on a chain of plaquettes, as shown in Fig. 1. The Kogut-Susskind Hamiltonian of the SU(2) lattice gauge theory is well-known [43]. In the electric basis [44, 45, 46], one can manage to project onto physical states that obey the Gauss law at each vertex and explicitly write down matrix elements of the Hamiltonian in the physical Hilbert space [47, 48, 49, 50]. One can then construct the physical Hamiltonian matrix on a lattice of given size with the local Hilbert space truncated at some electric flux value $j_{\rm max}$.
@@
We have multiplied the Hamiltonian by the lattice spacing $a$. So we will express energy and other physical quantities in units of $a$. For periodic boundary condition we set $\sigma^{z}_{i}=\sigma^{z}_{i+N}$ while for open boundary condition we choose $\sigma^{z}_{-1}=\sigma^{z}_{N}=-1$ (by this we mean the states at site $-1$ and $N$ are spin-down). We define the Hamiltonian density (per plaquette) as
@@ IV.1 Diffusive mode
We first calculate the symmetric correlator in frequency-momentum space on a periodic plaquette chain. Eigenstates can be identified by their eigenenergies and crystal momenta, $|E_{n}(p)\rangle$, since lattice translation operators commute with the Hamiltonian. The crystal momentum on a lattice of size $N$ is a multiple of $\frac{2\pi}{Na}$, i.e., $p=\frac{2\pi n_{p}}{N}$ in lattice units with $n_{p}\in[-\lfloor N/2\rfloor,-\lfloor N/2\rfloor+1,\cdots,\lfloor(N-1)/2\rfloor]$. If we treat lattice site labels $j$ and $i$ as continuous variables for the moment, inserting complete sets of eigenstates in Eq. (22) and Fourier transforming give
@@
On the lattice, we need to modify this expression in two aspects. Firstly, lattice positions are discrete so the spatial integration should be replaced with a summation over the lattice points. As a result, the Dirac delta function for momentum conservation turns to a Kronecker delta function
@@
Secondly, we have a finite number of eigenstates on the lattice and thus the energy spectrum is not continuous. We regularize the Dirac delta function for energy conservation with a rectangular function
@@ IV.2 Long-time tail
which is labeled by the red dashed line. The fitting is performed by using the results at $t>100$. The fitted parameter values are listed in Table 2. At the two couplings and two temperatures studied, $G_{s}(t,8,8)$ exhibits the power law decay at late time, which is consistent with the long-time tail of diffusive hydrodynamics. At the lower temperature $T=5$, some oscillating behavior is seen in $G_{s}(t,8,8)$, particularly in the $ag^{2}=0.8$ case. This is caused by the fact that at such a low temperature, not many states are active and thus the system is less quantum ergodic, as explained in the previous subsection. As a result, the time evolution phases of different active states are not canceled completely and the oscillation is still manifest.
@@ IV.3 Spatial diffusion
where $z=\frac{x^{2}}{4D_{\varepsilon}t}$, $S$ and $C$ denote the Fresnel integrals. Qualitatively, Eq. (IV.3) behaves similarly to the diffusion equation (17).
@@
The physical meaning of $c_{1}^{\prime}$ is $c_{1}^{\prime}\propto t^{-\frac{1}{2}}$ so we expect $\sqrt{t}c_{1}^{\prime}$ to be approximately independent of $t$ in the fit. This is approximately true as indicated in Table 3. The physical meaning of $c_{2}^{\prime}$ is $c_{2}^{\prime}=\frac{1}{4D_{\varepsilon}t}$ so we expect to be able to extract $D_{\varepsilon}=\frac{1}{4tc_{2}^{\prime}}$ from the fitted $c_{2}^{\prime}$. However, the values of $\frac{1}{4tc_{2}^{\prime}}$ as listed in Table 3 clearly show $t$ dependence. Furthermore, their magnitudes are one-order-of-magnitude smaller than the $D_{\varepsilon}$ values extracted from the transport peak in $G_{s}(\omega,k)$ as shown in Table 1. These two points indicate that using Eq. (IV.3) to fit $G_{x}(t,x)$ is not a suitable approach of extracting the energy diffusion coefficient in practice. The reason of the failure is that Eq. (IV.3) is obtained from the integral over $k$ as in Eq. (IV.3), which can differ significantly from the summation over $k=\frac{2\pi n_{k}}{N}$ on a small lattice. To demonstrate this claim, we fit $G_{s}(t,x)$ using the finitely summed version
@@ V.1 Quantum algorithm
    Thermal state preparation. The result of this step is to have the thermal density matrix, $\rho=\frac{1}{Z}e^{-\beta H}$ on the system qubits. We use the Quantum Imaginary Time Propagation (QITP) method [102, 103, 24, 104]. Details in the implementation for the SU(2) lattice gauge theory can be found in Ref. [24].
@@
    where $E_{T}$ is a constant and can be chosen to be the ground state energy of $H_{i}$, which is easy to obtain since $H_{i}$ is local and increases the success probability of the QITP algorithm. The non-unitary operator $e^{-\beta(H_{i}-E_{T})}$ is successfully applied on the system qubits if the ancilla state is measured in the |$\ket{0}$ state. We have applied this method in similar scenarios in our previous work [85], where more details can be found.
@@ V.2 Efficiency of the algorithm
$e^{-\frac{\beta}{2}H}=e^{-\frac{\beta}{2}H^{\rm el}}e^{-\frac{\beta}{2}H^{\rm mag}_{{\rm com}0}}e^{-\frac{\beta}{1}H^{\rm mag}_{{\rm com}2}}e^{-\frac{\beta}{2}H^{\rm mag}_{{\rm com}2}}\,,$ (39)
@@
The resources required in the QITP algorithm for implementing $e^{-\Gamma H_{i}}$ do not scale with the system volume since $H_{i}$ is the Hamiltonian density operator and thus local. Furthermore, we want $\Gamma$ to be small for the validity of Eq. (V.1), which corresponds to high temperature, justifying the QITP algorithm. Due to the same reason of locality, finding the unitary transformation that diagonalizes $H_{j}$ in the final measurement does not scale with the volume either. On the other hand, we expect resources for both steps to scale up with the cutoff $j_{\rm max}$. Since the size of the local Hilbert space grows polynomially with $j_{\rm max}$ roughly as $2j^{2}_{\rm max}$, and the minimum of $j_{\rm max}$ is inversely proportional to the accuracy of the calculation and the coupling [24], we expect the resources for the implementation of $e^{-\Gamma H_{i}}$ and the diagonalization of $H_{j}$ only scale polynomially with the final accuracy.
@@ V.3 Emulator results
We test the described quantum algorithm for an open plaquette chain of size $N=9$ with $ag^{2}=1$ and $j_{\rm max}=\frac{1}{2}$. Fig. 5 shows the obtained results of the symmetric correlators $G_{s}(t,j,i)$ for $i=j=4$ and $i=4,j=1$ at $\beta=0.05$, using a Trotterization time step of $\Delta t=0.05$ and $\Gamma=0.005$. The number of shots used in each case is $2\times 10^{6}$. Black solid line represents the exact evolution while the blue points represent the quantum circuit results obtained from the IBM emulator. We observe that the emulator results follow the exact evolution, up to statistical uncertainties. With the same number of shots, the uncertainty in the $i=4,j=1$ case looks relatively larger than the $i=j=4$ case, since the value of $G_{s}(t,1,4)$ is much smaller at early time. The uncertainty can be further reduced by using more shots. If we decrease the temperature $T$ in our calculations, more shots are also needed for the same accuracy because of the decreased success probability of thermal state preparation, as explained in the previous subsection.
@@ VI Conclusions
In this paper, we scrutinize emergent hydrodynamic behavior in real-time Hamiltonian dynamics of $2+1$-dimensional SU(2) lattice gauge theory on a plaquette chain, by computing symmetric correlation functions of stress-energy tensors. Due to the Umklapp processes on the lattice, momentum density is no longer an effective degree of freedom at high temperature and thus the hydrodynamics becomes energy diffusion. We found a transport peak near zero frequency in the symmetric correlator of energy densities. The width of the transport peak is approximately quadratically proportional to momentum at small momentum, when the lattice system is fully quantum ergodic. The transport peak leads to a power-law decay of the symmetric correlator at late time and can also explain the position dependence.
@@
In future work, we plan to investigate when the momentum density will become an effective degree of freedom again and thus the sound modes will become manifest. A necessary condition is the suppression of the Umklapp processes. As explained in Sec. II.2, the suppression happens when $T_{\rm phy}\ll\frac{\pi}{a}$. In physical units, we would like $T_{\rm phy}$ to be fixed. So in the continuum limit $a\to 0$, this condition will be satisfied. In lattice units, we want $T=aT_{\rm phy}\ll\pi$. However, the temperatures we used in this work are all above $5$, violating this condition. This is why we did not observe the sound modes. We cannot further lower the temperature in our current studies with $j_{\rm max}=\frac{1}{2}$, since the energy gap between the first excited state and the ground state is $E_{1}-E_{0}=4.1$ and $5.1$ for $ag^{2}=1.2$ and $ag^{2}=0.8$, respectively. Further lowering the temperature will only probe dynamics near the ground state rather than highly excited states in the deconfined region. As $a\to 0$, it is expected that $E_{1}-E_{0}\to 0$ in lattice units such that the physical value $\frac{E_{1}-E_{0}}{a}$ is finite and fixed. Then one is able to study the deconfined dynamics at a temperature satisfying $E_{1}-E_{0}\ll T\ll\pi$ in lattice units, with the Umklapp processes suppressed. For SU(2) pure gauge theory in $2+1$ dimensions, it is known that $ag^{2}\to 0$ as $a\to 0$ [112]. The fact that as we decrease $ag^{2}$ from 1.2 to 0.8, the energy gap $E_{1}-E_{0}$ increases, indicates that the finite-size artifacts in our studies with $j_{\rm max}=\frac{1}{2}$ are large and we have to increase the $j_{\rm max}$ cutoff. Currently, exact diagonalization can be done for $j_{\rm max}=1$ on an open $N=7$ chain [113] or a periodic $N=9$ chain [114]. However, these lattice sizes that allow exact diagonalization might be too small for the system to exhibit long-wavelength hydrodynamics. Other classical methods such as tensor networks or neutral networks should be explored, as well as quantum computing. It is unclear at the moment that if these classical methods can lead to the expected physics. If not, demonstrating the sounds modes for SU($N_{c}$) non-Abelian lattice gauge theory could be “quantum supremacy” for high-energy nuclear physics.

### v2

@@ Abstract
We search for emergent hydrodynamic modes in real-time Hamiltonian dynamics of $2+1$-dimensional SU(2) lattice gauge theory on a quasi one dimensional plaquette chain, by numerically computing symmetric correlation functions of energy densities on lattice sizes of about $20$ with the local Hilbert space truncated at $j_{\rm max}=\frac{1}{2}$. Because of the Umklapp processes, we only find a mode for energy diffusion. The symmetric correlator exhibits transport peak near zero frequency with a width approximately proportional to momentum squared at small momentum, when the system is fully quantum ergodic, as indicated by the eigenenergy level statistics. This transport peak leads to a power-law $t^{-\frac{1}{2}}$ decay of the symmetric correlator at late time, also known as the long-time tail, as well as diffusion-like spreading in position space. We also introduce a quantum algorithm for computing the symmetric correlator on a quantum computer and find it gives results consistent with exact diagonalization when tested on the IBM emulator. Finally we discuss the future prospect of searching for the sound modes.
@@ I Introduction
The values of shear viscosity and other transport coefficients of QCD at varying temperatures serve as important inputs for an array of phenomenological studies in heavy ion collisions. However, calculating them accurately in QCD is challenging due to nonperturbative effects and the sign problem in the Euclidean lattice QCD approach [13], although progress has been made [14, 15, 16, 17, 18, 19, 20, 21, 22]. This challenge together with the rapid development of quantum computing technology motivated recent studies on using the Hamiltonian formulation of lattice gauge theory to calculate the shear viscosity [23, 24]. The Hamiltonian lattice method is limited to only small systems at the moment [24], due to the limitation of current classical computers and quantum devices, which can lead to large systematic uncertainty. In particular, the lattice size may be too small to support the full development of long-wavelength hydrodynamic behavior. In this paper, we aim at answering this question and demonstrating one case where hydrodynamic behavior emerges in the Hamiltonian dynamics on the lattice. Because of the real-time nature, such a demonstration is difficult in the Euclidean lattice setup.
The main object that we will calculate in the Hamiltonian lattice approach is the real-time symmetric correlation function of stress-energy tensors, which is expected to evolve hydrodynamically at late time. Our results will show that a diffusive mode emerges in the SU(2) Hamiltonian dynamics on a lattice of size $\sim 20$. Because of the Umklapp effect, sound modes are not observed at the couplings we study with the local Hilbert space truncation.
@@ II.1 Hydrodynamics in the continuum
$\delta T^{\mu\nu}$ $=\begin{pmatrix}\delta\varepsilon_{0}&(\varepsilon_{0}+P_{0})\delta u^{x}\\ (\varepsilon_{0}+P_{0})\delta u^{x}&\delta P-\zeta\partial_{x}\delta u^{x}\end{pmatrix}\,.$ (4)
@@
$\partial_{t}\delta\varepsilon+\partial_{x}g^{x}=0\,,$
$\partial_{t}g^{x}+c_{s}^{2}\partial_{x}\delta\varepsilon-\gamma_{\zeta}\partial_{x}^{2}g^{x}=0\,,$ (6)
@@
$\partial_{t}^{2}g^{x}-c_{s}^{2}\partial_{x}^{2}g^{x}-\gamma_{\zeta}\partial_{t}\partial_{x}^{2}g^{x}=0\,.$ (7)
@@
$\omega_{s\pm}=\pm\sqrt{c_{s}^{2}k^{2}-\frac{\gamma_{\zeta}^{2}k^{4}}{4}}-i\frac{\gamma_{\zeta}k^{2}}{2}\,.$ (9)
@@ II.3 Real-time symmetric correlator
where $\rho_{T}\equiv\frac{e^{-\beta H}}{Z}$ with $\beta=1/T$ denotes the thermal density matrix, and $t_{0}$ and $x_{0}$ are some arbitrary reference points that can be set to zero. The symmetric correlator characterizes the fluctuations in the system at thermal equilibrium.
@@ III Hamiltonian of SU(2) Plaquette Chains
We will compute the symmetric correlator of the energy density for a specific lattice gauge theory, the SU(2) pure gauge theory in $2+1$ dimensions discretized on a chain of plaquettes, as shown in Fig. 1. The Kogut-Susskind Hamiltonian of the SU(2) lattice gauge theory is well-known [43]. In the electric basis [44, 45, 46], one can manage to project onto physical states that obey the Gauss law at each vertex and explicitly write down matrix elements of the Hamiltonian in the physical Hilbert space [47, 48, 49, 50]. One can then construct the physical Hamiltonian matrix on a lattice of a given size with the local Hilbert space truncated at some electric flux value $j_{\rm max}$.
@@
We have multiplied the Hamiltonian by the lattice spacing $a$. So we will express energy and other physical quantities in units of $a$. For the periodic boundary condition we set $\sigma^{z}_{i}=\sigma^{z}_{i+N}$ while for the open boundary condition we choose $\sigma^{z}_{-1}=\sigma^{z}_{N}=-1$ (by this we mean the states at site $-1$ and $N$ are spin-down). We define the Hamiltonian density (per plaquette) as
@@ IV.1 Diffusive mode
We first calculate the symmetric correlator in frequency-momentum space on a periodic plaquette chain. Eigenstates can be identified by their eigenenergies and crystal momenta, $|E_{n}(p)\rangle$, since lattice translation operators commute with the Hamiltonian. The crystal momentum on a lattice of size $N$ is a multiple of $\frac{2\pi}{Na}$, i.e., $p=\frac{2\pi n_{p}}{N}$ in lattice units with $n_{p}\in[-\lfloor N/2\rfloor,-\lfloor N/2\rfloor+1,\cdots,\lfloor(N-1)/2\rfloor]$. If we treat lattice site labels $j$ and $i$ as continuous variables for the moment, inserting complete sets of eigenstates in Eq. (22) and Fourier transforming gives
@@
On the lattice, we need to modify this expression in two aspects. First, lattice positions are discrete so the spatial integration should be replaced with a summation over the lattice points. As a result, the Dirac delta function for momentum conservation turns to a Kronecker delta function
@@
Second, we have a finite number of eigenstates on the lattice and thus the energy spectrum is not continuous. We regularize the Dirac delta function for energy conservation with a rectangular function
@@ IV.2 Long-time tail
which is labeled by the red dashed line. The fitting is performed by using the results at $t>100$. The fitted parameter values are listed in Table 2. At the two couplings and two temperatures studied, $G_{s}(t,8,8)$ exhibits the power-law decay at late time, which is consistent with the long-time tail of diffusive hydrodynamics. At the lower temperature $T=5$, some oscillating behavior is seen in $G_{s}(t,8,8)$, particularly in the $ag^{2}=0.8$ case. This is caused by the fact that at such a low temperature, not many states are active and thus the system is less quantum ergodic, as explained in the previous subsection. As a result, the time evolution phases of different active states are not canceled completely and the oscillation is still manifest.
@@ IV.3 Spatial diffusion
where $z=\frac{x^{2}}{4D_{\varepsilon}t}$, and $S$ and $C$ denote the Fresnel integrals. Qualitatively, Eq. (IV.3) behaves similarly to the diffusion equation (17).
@@
The physical meaning of $c_{1}^{\prime}$ is $c_{1}^{\prime}\propto t^{-\frac{1}{2}}$ so we expect $\sqrt{t}c_{1}^{\prime}$ to be approximately independent of $t$ in the fit. This is approximately true as indicated in Table 3. The physical meaning of $c_{2}^{\prime}$ is $c_{2}^{\prime}=\frac{1}{4D_{\varepsilon}t}$ so we expect to be able to extract $D_{\varepsilon}=\frac{1}{4tc_{2}^{\prime}}$ from the fitted $c_{2}^{\prime}$. However, the values of $\frac{1}{4tc_{2}^{\prime}}$ as listed in Table 3 clearly show $t$ dependence. Furthermore, their magnitudes are one-order-of-magnitude smaller than the $D_{\varepsilon}$ values extracted from the transport peak in $G_{s}(\omega,k)$ as shown in Table 1. These two points indicate that using Eq. (IV.3) to fit $G_{x}(t,x)$ is not a suitable approach of extracting the energy diffusion coefficient in practice. The reason for the failure is that Eq. (IV.3) is obtained from the integral over $k$ as in Eq. (IV.3), which can differ significantly from the summation over $k=\frac{2\pi n_{k}}{N}$ on a small lattice. To demonstrate this claim, we fit $G_{s}(t,x)$ using the finitely summed version
@@ V.1 Quantum algorithm
    Thermal state preparation. The result of this step is to have the thermal density matrix, $\rho_{T}=\frac{1}{Z}e^{-\beta H}$ on the system qubits. We use the Quantum Imaginary Time Propagation (QITP) method [102, 103, 24, 104]. Details in the implementation for the SU(2) lattice gauge theory can be found in Ref. [24].
@@
    where $E_{T}$ is a constant and can be chosen to be the ground state energy of $H_{i}$, which is easy to obtain since $H_{i}$ is local and increases the success probability of the QITP algorithm. The non-unitary operator $e^{-\Gamma(H_{i}-E_{T})}$ is successfully applied on the system qubits if the ancilla state is measured in the |$\ket{0}$ state. We have applied this method in similar scenarios in our previous work [85], where more details can be found.
@@ V.2 Efficiency of the algorithm
$e^{-\frac{\beta}{2}H}=e^{-\frac{\beta}{2}H^{\rm el}}e^{-\frac{\beta}{2}H^{\rm mag}_{{\rm com}0}}e^{-\frac{\beta}{2}H^{\rm mag}_{{\rm com}1}}e^{-\frac{\beta}{2}H^{\rm mag}_{{\rm com}2}}\,,$ (39)
@@
The resources required in the QITP algorithm for implementing $e^{-\Gamma H_{i}}$ do not scale with the system volume since $H_{i}$ is the Hamiltonian density operator and thus local. Furthermore, we want $\Gamma$ to be small for the validity of Eq. (V.1), which corresponds to high temperature, justifying the QITP algorithm. Because of the same reason of locality, finding the unitary transformation that diagonalizes $H_{j}$ in the final measurement does not scale with the volume either. On the other hand, we expect resources for both steps to scale up with the cutoff $j_{\rm max}$. Since the size of the local Hilbert space grows polynomially with $j_{\rm max}$ roughly as $2j^{2}_{\rm max}$, and the minimum of $j_{\rm max}$ is inversely proportional to the accuracy of the calculation and the coupling [24], we expect that the resources for the implementation of $e^{-\Gamma H_{i}}$ and the diagonalization of $H_{j}$ will only scale polynomially with the final accuracy.
@@ V.3 Emulator results
We test the described quantum algorithm for an open plaquette chain of size $N=9$ with $ag^{2}=1$ and $j_{\rm max}=\frac{1}{2}$. Figure 5 shows the obtained results of the symmetric correlators $G_{s}(t,j,i)$ for $i=j=4$ and $i=4,j=1$ at $\beta=0.05$, using a Trotterization time step of $\Delta t=0.05$ and $\Gamma=0.005$. The number of shots used in each case is $2\times 10^{6}$. Black solid line represents the exact evolution while the blue points represent the quantum circuit results obtained from the IBM emulator. We observe that the emulator results follow the exact evolution, up to statistical uncertainties. With the same number of shots, the uncertainty in the $i=4,j=1$ case looks relatively larger than the $i=j=4$ case, since the value of $G_{s}(t,1,4)$ is much smaller at early time. The uncertainty can be further reduced by using more shots. If we decrease the temperature $T$ in our calculations, more shots are also needed for the same accuracy because of the decreased success probability of thermal state preparation, as explained in the previous subsection.
@@ VI Conclusions
In this paper, we scrutinize emergent hydrodynamic behavior in real-time Hamiltonian dynamics of $2+1$-dimensional SU(2) lattice gauge theory on a plaquette chain, by computing symmetric correlation functions of stress-energy tensors. Because of the Umklapp processes on the lattice, momentum density is no longer an effective degree of freedom at high temperature and thus the hydrodynamics becomes energy diffusion. We found a transport peak near zero frequency in the symmetric correlator of energy densities. The width of the transport peak is approximately quadratically proportional to momentum at small momentum when the lattice system is fully quantum ergodic. The transport peak leads to a power-law decay of the symmetric correlator at late time and can also explain the position dependence.
@@
In future work, we plan to investigate when the momentum density will become an effective degree of freedom again and thus the sound modes will become manifest. A necessary condition is the suppression of the Umklapp processes. As explained in Sec. II.2, the suppression happens when $T_{\rm phy}\ll\frac{\pi}{a}$. In physical units, we would like $T_{\rm phy}$ to be fixed. So in the continuum limit $a\to 0$, this condition will be satisfied. In lattice units, we want $T=aT_{\rm phy}\ll\pi$. However, the temperatures we used in this work are all above $5$, violating this condition. This is why we did not observe the sound modes. We cannot further lower the temperature in our current studies with $j_{\rm max}=\frac{1}{2}$, since the energy gap between the first excited state and the ground state is $E_{1}-E_{0}=4.1$ and $5.1$ for $ag^{2}=1.2$ and $ag^{2}=0.8$, respectively. Further lowering the temperature will only probe dynamics near the ground state rather than highly excited states in the deconfined region. As $a\to 0$, it is expected that $E_{1}-E_{0}\to 0$ in lattice units such that the physical value $\frac{E_{1}-E_{0}}{a}$ is finite and fixed. Then one is able to study the deconfined dynamics at a temperature satisfying $E_{1}-E_{0}\ll T\ll\pi$ in lattice units, with the Umklapp processes suppressed. For SU(2) pure gauge theory in $2+1$ dimensions, it is known that $ag^{2}\to 0$ as $a\to 0$ [112]. The fact that as we decrease $ag^{2}$ from 1.2 to 0.8, the energy gap $E_{1}-E_{0}$ increases, indicates that the finite-size artifacts in our studies with $j_{\rm max}=\frac{1}{2}$ are large and we have to increase the $j_{\rm max}$ cutoff. Currently, exact diagonalization can be done for $j_{\rm max}=1$ on an open $N=7$ chain [113] or a periodic $N=9$ chain [114]. However, these lattice sizes that allow exact diagonalization might be too small for the system to exhibit long-wavelength hydrodynamics. Other classical methods such as tensor networks or neutral networks should be explored, as well as quantum computing. It is unclear at the moment if these classical methods can lead to the expected physics. If not, demonstrating the sounds modes for SU($N_{c}$) non-Abelian lattice gauge theory could be “quantum supremacy” for high-energy nuclear physics.